Accessibility Assistance

Skip to Content




Information Technology
George A. Smathers Libraries
University of Florida
Suite 1250
2124 NE Waldo Rd
Gainesville, FL 32609

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
January 11, 2011

Present: Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Winston Harris, Michael Jay, Maureen Kelly, Tom Minton, Chris Nicolich, Amy Polk, Patricia Ruwell, Josh Spurgin, Mark Sullivan, Mike Tyler, Evan Wack
Announcements
  1. MS Academy link: http://www.cns.ufl.edu/training/
  2. New UF VPN announced: http://net-services.ufl.edu/provided_services/vpn
  3. Review of the meeting minutes from the IT Directors.
  4. Email regarding the preservation of all data related to donors/grants - do not delete.
Review Minutes of 1/4/2011
  1. Announcement #1 removed.
  2. HW#8 - "A proposal for laptop/classroom replacement of technology is being developed by Diane Bruxvoort" added.
  3. Ntwk/Server #1 - "Meeting may have been moved to January 13 due to ALA" added.
Ongoing IT Updates

Hardware
  1. Altiris Notification Server - online, configuring plug-in.
  2. Altiris Deployment Server - bring it online. Once the Deployment Server is up, it can be taken off the virtual server.
  3. AV proposals in the works: Grand Reading Room, Group Study Rooms and Education.
  4. SharePoint - Debra is getting back to work it.
  5. Public WIN7 migration - pre-staging is 90% complete. Logan to work on PXE boot. Awaiting new Clean Slate version. Updating image. MSL has volunteered to use MSL308 as a test bed.
  6. Remedy - no decision.
  7. MSL 107 technology proposal - awaiting word.
  8. Help Desk - still need to hire another student.
  9. Tapes: 1769 total. 116 SPOT to be stored at ALF until they can be converted and 16 Chiles to be checked out to John Nemmers in Special Collections. He will ship them out for conversion. If the conversion is successful, hopefully they can be destroyed by the same company.
Software
  1. Bad data in source files for Patron Updates - ongoing.
  2. Recruitment Study (HR) - Bonnie and Brian are out at ALA.
  3. Document disposition database-User management tool - Tom is working on it. Will probably be another week or two. It is a tool that allows new account set up.
  4. Document disposition database - Phase I launched Jan. 3. Working on punch list items (improvements).
  5. Macros - donation book plate revision almost done. Macro for Betsy is pending.
  6. Kuali OLE - coding ramp up in Indiana. Maureen is a rep on the technical council. They are putting together specs.
  7. ILL server switch - scheduled for Thursday of next week. NT 18 is done.
  8. Panama Canal file migration - lots of action lately. They are exporting a list for Chris to turn into a useable file. Rachel and LAC are involved.
  9. Library Hours - Shruti has been out.
  10. Library class registration system (MSL) - Maureen is looking into the HSC software to see if we can use it.
  11. Integrating "Send to Outlook" into ITS - Chris is working on it.
Network/Server Administration
  1. ALF network issues - still troubleshooting and trying to get a firm date for fiber. Ticket for networking build-out for work trailers at ALF submitted.
  2. SharePoint - Amy is meeting with Branch Chairs on January 26 to discuss their site. We are installing SharePoint Designer template and tutorials to get a handle on branding.
  3. DHCP move in process; hope to begin switching to Library server next week. Will also fix PXE pointer next week. Discussion on ALEPH Circulation authorization and DHCP roll out.
  4. Exchange issues resolved. SharePoint still asks for log in.
  5. NT13 - talked to Joe - some issues remain; may need to virtualize it locally depending on time frame for fix.
  6. West Security - will be scheduled at the same time as Vmware rack move. The rest of the hardware is almost ready for survey/delivery to other departments.
ILS
  1. Starting to pick back up again.
  2. Betsy has suggested to Cynthia to use the Data Warehouse tables. That data yields different results. Matthew Daley can help with this issue.
DLC General Issues
  1. Still working to merge traffic into the UFDC web site.
  2. Randall has purchased processor upgrades for 5 of the DLC Copibooks. The Hardware Unit will need to install. They have requested we talk to Gus, Randall or Traveler prior to chip installation.
Web Projects
  1. Study Room Reservation - waiting for feedback, made one change. Getting standalone machine ready. Will use Squid software to lock it down. Will be ready next week. This will give patrons one machine to use to reserve the Study Rooms.
  2. Academic and Scholarly Outreach web site - Logan migrated the template. Awaiting content from Dr. Silver. The navigation works.
  3. Serial Solutions stopped organizing by subject. People were noticing no returns on their searches. Will contact the vendor to see if the change is permanent. It is an issue when vendors change their sites and don't let us know.
  4. Virtual Bookmarks - images uploaded and java script edited. There is a new image for Judaica. It was not working right. The name was not lining up properly.
  5. Logan notes that the Assignment Calculator runs on an old server that was shut down. He will put it up on a new box. Jabber, as well.